# SADDE: Semi-supervised Anomaly Detection with Dependable Explanations

### v1

@@ 2.2. Global-local Knowledge Association Mechanism (KAM)
Based on the results obtained by the global interpreter $G$ and local interpreter $L$, a feature similarity calculation method (as defined in Eq.(1)) is used to further verify the reliability of the interpretation results. Inspired by (Fan et al., 2020), the similarity score is defined as follows. Different from (Fan et al., 2020), it measures the similarity between the locally interpreted key feature set of an instance and the globally interpreted key feature set of a cluster that the instance belongs to.
(1) $sim(int_{x}(m),int_{y}(m^{\prime}),k)=2*\frac{int^{k}_{x}(m)\cap int^{k}_{y}(m^{\prime})}{\left|int^{k}_{x}(m)\right|+\left|int^{k}_{y}(m^{\prime})\right|},$
@@
The local interpretations of an instance are considered highly reliable when the feature similarity of the local and global interpretations is higher than a given Threshold. In our case study, we utilize the DeepLift (Shrikumar et al., 2017b) as our local interpreter and PFIE (Fisher et al., 2019) as our global interpreter as an example. Other local/model-agnostic or global interpreters can also be used as the local and global interpreters in KAM. Note that different from existing interpreters such as (Han et al., 2021; Liu et al., 2017a; Yang et al., 2021), which only interpret anomalous samples or outliers, SADDE interprets both normal and anomalous samples for that both normal and anomalous samples can potentially be used to update the model in SADDE, thereby influencing the entire learning and decision-making process of anomaly detection. Besides, in security applications, a transparent decision-making process is crucial for system operators of security applications to adopt theoretical algorithms in practice.
2.3. Two-stage Semi-supervised Learning (ToS)
@@
Double Verification Mechanism. The goal of the double verify mechanism is to further reduce false positive and miss-detection rates and improve detection accuracy through a two-layer verification process that incorporates both the similar feature verification in KAM and the global and local prediction check in ToS. For the first-layer verification, as aforementioned in Section 2.2, it compares the calculated similarity score (see Eq.(1)), denoted as $sim$, with a predefined threshold $T$. We consider the interpretations to be reliable if $sim>T$. To some extent, it indicates that the sample has not been modified or attacked and could be used for the next round of re-training. For the second-layer verification, we consider the generated pseudo labels to be confident when the local model’s prediction $\hat{y}_{l}$ matches the global model’s prediction $\hat{y}_{g}$. The selected reliable samples and their pseudo-labels are used to retrain both the global and clustering models. The model then waits for new network flow data to arrive for further processing.
In our case study, we use the Multilayer Perceptron Network (MLP) (Rumelhart et al., 1986) and DBSCAN (Ester et al., 1996) as our local and global models respectively as an example for evaluating the effectiveness of the proposed ToS. Other machine learning models can also be chosen accordingly for specific tasks.
3. Evaluation
@@
This section presents the experimental setup and results. The datasets are detailed in subsection 3.1. Next, the experiment setup is introduced in subsection 3.2. Finally, in subsection 3.3, we compare the proposed KAM and ToS with five and eight state-of-the-art on two representative network anomaly detection datasets and analyze their results.
3.1. Dataset
@@
Two widely used network anomaly detection datasets, ISCXTor2016 (Lashkari et al., 2017) and CIC-DoHBrw-2020 (MontazeriShatoori et al., 2020), are employed to evaluate the performance of SADDE. ISCXTor2016 contains Tor and non-Tor network flows in various applications and each instance has 28 features. It consists of 68,191 samples, of which 56,534 are non-Tor and 11,657 are Tor samples. CIC-DoHBrw-2020 contains labeled traffic generated by DNS over HTTPS (DoH) protocols. It is a dataset for detecting browser-based DoH activities and studying secure communication anomalies. Each instance has 33 features. There are 933,189 instances, including 216,649 DoH and 716,540 non-DoH samples.
@@
For the ISCXTor2016 dataset, the first 1% of the data is used as the pre-training set (i.e., small labeled set), the last 20% as the testing set, and the next 20% as the validation set, while the remaining data is allocated for re-training. The CSIC-DoHBrw-2020 dataset follows a similar partitioning structure, except that the first 0.5% of the data is used as the pre-training set.
3.2. Setup
@@
System Settings. The experiments are conducted using Python (v3.8), along with the PyTorch framework (v1.11.0), CUDA (v11.3), and Scikit-learn (v1.3.2) to develop and evaluate SADDE. A three-layer FCN is employed as our $M_{1}$, and each layer contains 256 neurons.
@@
During training, the Adam optimizer is used with a learning rate of $31e-$ 3. A dropout rate of 0.3 is implemented to avoid over-fitting, and cross-entropy loss was selected as the objective function for classification tasks. We set the batch size to 64. DBSCAN was employed with an epsilon value of 0.5 and a minimum sample size of 1 for clustering, ensuring that all samples could potentially be included in a cluster. The output of FCN’s third fully connected layer is utilized as the input of DBSCAN. Additionally, early stopping was implemented with a patience of 5 epochs during both the pre-training and re-training phases so that the training process could be halted if there is no improvement in the validation loss over 5 epochs. In PFIE, each feature is permuted by 100 times. For DeepLift, the eps parameter is configured to $101e-$ 10. The number of key features is specified as 10, and the threshold for the similarity score is established at 0.6.
@@
Benchmarks. We employ a comprehensive set of benchmarks in this work to validate the performance of SADDE. Firstly, we conduct a comparative analysis of the proposed KAM against eight state-of-the-art interpreters, including Lime (Ribeiro et al., 2016b), SHAP (Lundberg and Lee, 2017), DeepLift (Shrikumar et al., 2017b), COIN (Liu et al., 2017b), DiCE (Mothilal et al., 2020), xNIDS (Wei et al., 2023), CADE (Yang et al., 2021), Anchors (Ribeiro et al., 2018).
@@
Then, we compared the introduced ToS with five state-of-the-art semi-supervised methods, including InstantT (Li et al., 2024), ACR (Wei and Gan, 2023), Clustering* (Fini et al., 2023), LVM (Min et al., 2024), and M3S (Sun et al., 2020a).
@@
Evaluation Metrics. We use Accuracy (Acc.), Precision (Pre.), Recall (Rec.), F1-score (F1.), False Alarm Rate (FAR), Missed Detection Rate (MDR), and standardized partial AUC (SPAUC) (McClish, 1989) to measure different semi-supervised models’ performance. Particularly, SPAUC measures the performance of a model within a specific region of the ROC curve, focusing on a segment that is most relevant under conditions of class imbalance and differential costs of misclassification. FAR reflects the amount of normal network traffic that is wrongly classified as abnormal while MDR indicates how much anomaly network traffic that are miss-detected by the models. The higher the SPAUC, the better the performance is. The lower the FAR and MDR, the better the performance is. Here we use $\textit{SPAUC}_{\textit{FPR}\leq 0.05}$ for all experiments.
@@
Figure 3. Fidelity, stability, and robustness evaluation of interpreters on ISCXTor2016.
Figure 4. Fidelity, stability, and robustness evaluation of interpreters on CIC-DoHBrw-2020.
  • Fidelity. To evaluate the fidelity of an interpreter, we define an indicator similar to that used in (Han et al., 2021), called the Label Flipping Rate (LFR), which is the ratio of samples that change to a different class after being modified based on the interpretation results. Specifically, in the context of a classification task, the interpreter identifies several key features. The values of these key features are replaced with the average values corresponding to the opposite class, and a new prediction is generated to determine whether the label has been changed. The LFR is mathematically expressed as:
    (3) $LFR=\frac{\sum_{i=1}^{n}I(f^{\prime}(x_{i})\neq f(x_{i}))}{n}$
@@
    (4) $Stability=\frac{1}{n}\sum_{i=1}^{n}\text{sim}(int_{x}(m)_{1}^{i},int_{x}(m)_{2}^{i})$
@@
    (5) $Robustness=\frac{1}{n}\sum_{i=1}^{n}\text{sim}(int_{x}(m)_{i},int_{x}(m+\epsilon)_{i},k)$
@@
3.3. Results and Analysis
Verification of Model Pre-training. Table 1 and Table 2 present the performance of ToS when trained on the pre-training set (denoted as ①) and when trained on both the pre-training and re-training sets (denoted as ②). The results indicate that the re-training approach (②) achieves significantly better performance than pre-training alone (①), with improvements of approximately 20%, 10%, and 14% in Precision, FAR, and SPAUC, on the ISCXTor2016 dataset. Similarly, on the CIC-DoHBrw-2020 dataset, the re-training process (②) enhances the model’s performance by about 14%, 26%, and 4% in terms of F1-score, MDR, and SPAUC, compared to training only on the smaller pre-training set (①). These findings demonstrate the effectiveness of the semi-supervised learning process in improving the performance of the proposed SADDE.
Table 1. Performance (%) of ToS on ISCXTor2016: comparisons between model performance (%) with only the pre-training set (①) and with both pre-training and re-training set (②). $\triangledown$ denotes the performance difference relative to the former. The best metric performance is bolded.
@@ 3.3. Results and Analysis
Table 2. Performance (%) of ToS on CIC-DoHBrw-2020: comparisons between model performance (%) with only the pre-training set (①) and with both pre-training and re-training set (②). $\triangledown$ denotes the performance difference relative to the former. The best metric performance is bolded.
@@
Figure 5. Efficiency evaluation of interpreters.
Figure 6. Performance of ToS compared with the state-of-the-art semi-supervised models: F1-score vs. SPAUC.
Figure 7. Performance of ToS compared with the state-of-the-art semi-supervised models: FAR vs. MDR.
Comparison with the State-of-the-art Interpreters. Figures 3 and 4 illustrate the evaluation of various interpreters across the ISCXTor2016 and CIC-DoHBrw-2020 datasets, highlighting performance differences across three critical dimensions: fidelity, stability, and robustness. Fig. 5, in turn, presents the results of the efficiency evaluation. KAM demonstrates a superior balance across all these dimensions compared to other methods. Notably, while COIN outperforms KAM in terms of fidelity and robustness on the ISCXTor2016 dataset, its efficiency is significantly lower, being nearly 100 times slower than KAM.
@@
We observe that in the ISCXTor2016 dataset, the fidelity of all interpreters is relatively low. This is because anomaly samples, such as Tor traffic, are more susceptible to label flipping, as modifications to their key features are more likely to alter model predictions. However, the dataset’s significant class imbalance—reflected in a normal-to-anomalous ratio of approximately 4.85:1—leads to a lower overall LFR. In contrast, the CIC-DoHBrw-2020 dataset has a normal-to-anomalous ratio of about 3.31:1, offering a relatively more balanced scenario for evaluating fidelity.
According to Fig. 5, KAM demonstrates a balanced performance between interpretability, reliability, and efficiency. While DeepLift and xNIDS are more efficient compared to our method, they do not perform as well on other key metrics. Moreover, we observed that Anchors is less efficient on the CIC-DoHBrw-2020 dataset compared to the ISCXTor2016 dataset. This is due to the combinatorial optimization involved in the search for an anchor, which causes a significant drop in efficiency as the number of features increases. With 28 features in the ISCXTor2016 dataset and 33 in the CIC-DoHBrw-2020 dataset, this increase in feature count directly contributes to the observed difference in performance.
@@
Comparison with the State-of-the-art Semi-supervised Models. As shown in Table 3 and Table 4, ToS outperforms other methods. ToS achieves SPAUC as high as 67.42% and 55.94% on ISCXTor2016 and CIC-DoHBrw-2020 datasets with about 1.33% and 1.53% improvement, respectively. Notably, LVM performs similarly to ToS on the ISCXTor2016 dataset, with only a 0.82% difference in f1-score and 1.33% difference in SPAUC. However, its performance on the CIC-DoHBrw-2020 dataset is significantly lower than ToS, with a 19.47% gap in f1-score and 4.48% in SPAUC, indicating that ToS exhibits better stability and is more adaptable to different scenarios.
@@
Furthermore, as illustrated in Fig. 7, ToS strikes the optimal balance between FAR and MDR. Although the Clustering* method exhibits an exceptionally low FAR on the ISCXTor2016 dataset, its MDR is notably high, nearing 100%. Similarly, while instantT demonstrates superior FAR performance compared to ToS on the CIC-DoHBrw-2020 dataset, both suffer from significantly elevated MDR values. As shown in Fig. 6, ToS also excels in balancing F1-score and SPAUC, outperforming other models on both two datasets.
@@
Table 3. Performance (%) of ToS compared with the state-of-the-art semi-supervised models on ISCXTor2016.
@@
Table 4. Comparison with the state-of-the-art semi-supervised models’ performance (%) on CIC-DoHBrw-2020.
@@
4. Discussion
Although in this work we only evaluated SADDE on network anomaly detection tasks as an example, the proposed KAM and ToS can be utilized in a wide range of semi-supervised anomaly detection applications. Besides, the robustness evaluation of AnoamlyIAP is mainly considered against data poisoning/feature perturbation attacks (See Section 2.2). There are other attacks that may fail SADDE like model poisoning attacks because SADDE is highly dependent on the local and global models’ predictions. Future work can investigate the SADDE’s robustness against more attacks. Moreover, hyper-parameters of SADDE are configured empirically. Future research can design more advanced approaches to configure hyperparameters. In this work, different data distribution imbalance ratio is not experimented due to the page limitation. Research on the influence of different imbalance ratios or the performance of SADDE under extreme data distributions can be conducted in the future. Furthermore, more anomaly detection scenarios with different model combinations can also be tested in the future to explore the generalization performance of SADDE. Finally, it is worthy of exploring how to use SADDE on scenarios with only a few labeled samples, for example, by developing better multi-stage semi-supervised learning approaches or new matching strategies.
5. Related Work

### v4

@@ 4.2 Global-local Knowledge Association Mechanism (KAM)
Based on the results obtained by the global interpreter $G$ and local interpreter $L$, a feature similarity calculation method (as defined in Eq.(5)) is used to further verify the reliability of the interpretation results. Following [fan2020can], the similarity score is defined as follows. It measures the similarity between the locally interpreted key feature set of an instance and the globally interpreted key feature set of a cluster that the instance belongs to.
$sim(int_{x}(m),int_{y}(m^{\prime}),k)=2*\frac{int^{k}_{x}(m)\cap int^{k}_{y}(m^{\prime})}{\left|int^{k}_{x}(m)\right|+\left|int^{k}_{y}(m^{\prime})\right|},$ (5)
@@
The local interpretations of an instance are considered highly reliable when the feature similarity of the local and global interpretations is higher than a given Threshold. In our case study, we utilize the DeepLift [shrikumar2017learning_deeplift] as our local interpreter and PFIE [fisher2019all] as our global interpreter as an example. Other local/model-agnostic or global interpreters can also be used as the local and global interpreters in KAM. Note that different from existing interpreters such as [han2021deepaid, liu2017contextual_coin, yang2021cade], which only interpret anomalous samples or outliers, AnomalyAID interprets both normal and anomalous samples for that both normal and anomalous samples can potentially be used to update the model in AnomalyAID, thereby influencing the entire learning and decision-making process of anomaly detection. Besides, in security applications, a transparent decision-making process is crucial for system operators of security applications to adopt theoretical algorithms in practice.
4.3 Two-stage Semi-supervised Learning (ToS)
@@
Double Verification Mechanism. The goal of the double verify mechanism is to further reduce false positive and miss-detection rates and improve detection accuracy through a two-layer verification process that incorporates similar feature verification in KAM and global and local prediction checks in ToS. For the first-layer verification, as aforementioned in Section 4.2, it compares the calculated similarity score (see Eq.(5)), denoted as simsim, with a predefined threshold $T$. We consider the interpretations to be reliable if $sim>T$. To some extent, it indicates that the sample has not been modified or attacked and could be used for the next round of re-training. For the second-layer verification, we consider the generated pseudo labels to be confident when the local model’s prediction $\hat{y}_{l}$ matches the global model’s prediction $\hat{y}_{g}$. The selected reliable samples and their pseudo-labels are used to retrain both the global and clustering models. The model then waits for new network flow data to arrive for further processing.
5 Evaluation
@@
This section presents the experimental setup and results. The datasets are detailed in subsection 5.1. Next, the experiment setup is introduced in subsection 5.2. Finally, in subsection 5.3, we compare the proposed KAM and ToS with 13 state-of-the-art algorithms in three representative network anomaly detection datasets and analyze their results.
5.1 Dataset
Table 1: Descriptions of three widely-used network anomaly detection datasets used in our experiments.
Dataset Pre-training Training Validation Testing
#Malicious #Benign #Malicious #Benign #Malicious #Benign #Malicious #Benign
ISCXTor2016 691 150 41735 8777 13146 2850 14108 2730
CIC-DoHBrw-2020 4446 1350 534045 161499 173402 52650 178049 53800
UNSW-NB15 560 1193 44352 94518 11088 23630 37000 45332
@@ 5.1 Dataset
Three widely used network anomaly detection datasets, ISCXTor2016 [lashkari2017characterization], CIC-DoHBrw-2020 [montazerishatoori2020detection], and UNSW-NB15 [moustafa2015unsw], are employed to evaluate the performance of AnomalyAID. ISCXTor2016 contains Tor and non-Tor network flows in various applications. Each instance has 28 features, including Tor and non-Tor flows across various applications, supporting privacy and security research. It consists of 68,191 samples, of which 56,534 are non-Tor and 11,657 are Tor samples. CIC-DoHBrw-2020 contains labeled traffic generated by DNS over HTTPS (DoH) protocols. It is a dataset for detecting browser-based DoH activities and studying secure communication anomalies. Each instance has 33 features. There are 933,189 instances, including 216,649 DoH and 716,540 non-DoH samples. The UNSW-NB15 dataset includes 100 GB of network traffic captured using tcpdump, containing nine types of attacks and 43 features generated by 12 algorithms. It is designed to support security research on modern network activities and attacks. There are 257,673 instances, including 93,000 benign and 164,673 attack samples.
@@
For the ISCXTor2016 dataset and UNSW-NB15 dataset, the first 1% of the data is used as the pre-training set (i.e., small labeled set), the last 20% as the testing set, and the next 20% as the validation set, while the remaining data is allocated for re-training. The CSIC-DoHBrw-2020 dataset follows a similar partitioning structure, except that the first 0.5% of the data is used as the pre-training set. Detailed data distributions of the three datasets for pre-training, training, validation, and testing are presented in Table 1.
Note that although in this work we only evaluated AnomalyAID on network anomaly detection tasks as an example, the proposed KAM and ToS can be utilized in a wide range of semi-supervised anomaly detection applications.
5.2 Setup
@@
System Settings. The experiments are conducted using Python (v3.8), along with the PyTorch framework (v1.11.0), CUDA (v11.3), and Scikit-learn (v1.3.2) to develop and evaluate AnomalyAID. A three-layer FCN is employed as our $M_{1}$, and each layer contains 256 neurons.
@@
During training, the Adam optimizer is used with a learning rate of 1e31e-3. A dropout rate of 0.3 is implemented to avoid over-fitting, and cross-entropy loss was selected as the objective function for classification tasks. We set the batch size to 64. DBSCAN was employed with an epsilon value of 0.5 and a minimum sample size of 1 for clustering, ensuring that all samples could potentially be included in a cluster. The output of FCN’s third fully connected layer is utilized as the input of DBSCAN. Additionally, early stopping was implemented with a patience of 5 epochs during both the pre-training and re-training phases so that the training process could be halted if there was no improvement in the validation loss over 5 epochs. In PFIE, each feature is permuted 100 times. For DeepLift, the eps parameter is configured to 1e101e-10. The number of key features is specified as 10, and the threshold for the similarity score is established at 0.6 from extensive experiments.
@@
Benchmarks. We employ a comprehensive set of benchmarks in this work to validate the performance of AnomalyAID. Firstly, we conduct a comparative analysis of the proposed KAM against eight state-of-the-art interpreters, including Lime [ribeiro2016should_lime], SHAP [lundberg2017unified_shap], DeepLift [shrikumar2017learning_deeplift], COIN [liu2017contextual_coin], DiCE [mothilal2020explaining], xNIDS [wei2023xnids], CADE [yang2021cade], Anchors [ribeiro2018anchors], and CETP [lin2024cetp].
  • 1.
@@
  • 9.
    DeepSHAP: It combines DeepLIFT and Shapley values to compute feature attribution.
Then, we compared the introduced ToS with five state-of-the-art semi-supervised methods, including InstantT [li2024instantT], ACR [wei2023towards_ACR], Clustering* [Fini_2023_CVPR_clustering], LVM [min2024leveraging_LVM], and M3S [sun2020multi_M3S].
  • 1.
@@
  • 6.
    CETP: It assigns pseudo-labels to unlabeled data through confidence thresholds and dynamic loss weighting strategies, thereby optimizing classification performance.
Evaluation Metrics. We use Accuracy (Acc.), Precision (Pre.), Recall (Rec.), F1-score (F1.), False Alarm Rate (FAR), Missed Detection Rate (MDR), and standardized partial AUC (SPAUC)to measure different semi-supervised models’ performance. Particularly, SPAUC measures the performance of a model within a specific region of the ROC curve, focusing on a segment that is most relevant under conditions of class imbalance and differential costs of misclassification. FAR reflects the amount of normal network traffic that is wrongly classified as abnormal, while MDR indicates how much anomaly network traffic that are miss-detected by the models. The higher the SPAUC, the better the performance is. The lower the FAR and MDR, the better the performance is. Here we use $\textit{SPAUC}_{\textit{FPR}\leq 0.05}$ for all experiments.
@@
  • 1.
    Fidelity. To evaluate the fidelity of an interpreter, we define an indicator similar to that used in [han2021deepaid], called the Label Flipping Rate (LFR), which is the ratio of samples that change to a different class after being modified based on the interpretation results. Specifically, in the context of a classification task, the interpreter identifies several key features. The values of these key features are replaced with the average values corresponding to the opposite class, and a new prediction is generated to determine whether the label has been changed. The LFR is mathematically expressed as:
    $\textit{LFR}=\frac{\sum_{i=1}^{n}I(f^{\prime}(x_{i})\neq f(x_{i}))}{n}$ (7)
@@
    $Stability=\frac{1}{n}\sum_{i=1}^{n}\text{sim}(int_{x}(m)_{1}^{i},int_{x}(m)_{2}^{i})$ (8)
@@
    $Robustness=\frac{1}{n}\sum_{i=1}^{n}\text{sim}(int_{x}(m)_{i},int_{x}(m+\epsilon)_{i},k)$ (9)
@@
5.3 Results and Analysis
Table 2: Performance (%) of ToS on ISCXTor2016: comparisons between model performance (%) with only the pre-training set (①) and with both pre-training and re-training set (②). $\triangledown$ denotes the performance difference relative to the former. The best metric performance is bolded.
@@ 5.3 Results and Analysis
Table 3: Performance (%) of ToS on CIC-DoHBrw-2020: comparisons between model performance (%) with only the pre-training set (①) and with both pre-training and re-training set (②). $\triangledown$ denotes the performance difference relative to the former. The best metric performance is bolded.
@@
Table 4: Performance (%) of ToS on UNSW-NB15: comparisons between model performance (%) with only the pre-training set (①) and with both pre-training and re-training set (②). $\triangledown$ denotes the performance difference relative to the former. The best metric performance is bolded.
Setting Acc. Pre. Rec. F1. FAR MDR SPAUC
72.93 72.65 72.56 72.60 31.17 23.70 51.85
78.38 79.38 78.38 78.38 10.73 30.49 57.02
$\triangledown$ +5.45 +6.73 +5.82 +5.78 -20.44 +6.79 +5.17
Figure 4: Fidelity, stability, and robustness evaluation of interpreters on ISCXTor2016.
Figure 5: Fidelity, stability, and robustness evaluation of interpreters on CIC-DoHBrw-2020.
Figure 6: Fidelity, stability, and robustness evaluation of interpreters on UNSW-NB15.
5.3.1 Verification of Model Pre-training.
Table 2, Table 3, and Table 4 present the performance of ToS when trained on the pre-training set (denoted as ①) and when trained on both the pre-training and re-training sets (denoted as ②). The results indicate that the re-training approach (②) achieves significantly better performance than pre-training alone (①) on most of the evaluation metrics in the ISCXTor2016, CIC-DoHBrw-2020, and UNSW-NB15 datasets. Specifically, it achieves 67.42%, 55.94%, and 57.02% SPAUC on the ISCXTor2016, CIC-DoHBrw-2020, and UNSW-NB15 dataset, with improvements of approximately 13.92%, 3.60%, and 5.17% in SPAUC compared with ①. In particular, the performance of ToS on setting ② is 20%, 10%, and 14% better in Precision, FAR, and SPAUC, on the ISCXTor2016 dataset (as presented in Table 2). Similarly, on the CIC-DoHBrw-2020 dataset, the re-training process (②) enhances the model’s performance by about 14%, 26%, and 4% in terms of F1-score, MDR, and SPAUC, compared to training only on the smaller pre-training set (①). Additionally, as shown in Table 4, the re-training process (②) improves the model’s performance by approximately 5.45% in accuracy, 6.73% in precision, 5.82% in F1-score, 5.78% in recall, 20.44% in FAR, and enhances the SPAUC by 5.17%, compared to training exclusively on the pre-training set (①) in the UNSW-NB15 dataset. Overall, ToS performs better when trained on both the pre-training and re-training sets (denoted as ②) compared to when it only trained on the pre-training set (denoted as ①), despite that the MDR of the latter (①) is 0.44% and 6.79% worse than the former (②) on the ISCXTor2016 and UNSW-NB15 datasets. These findings demonstrate the effectiveness of the semi-supervised learning process (i.e., ToS) in improving the performance of the proposed AnomalyAID.
Figure 7: Efficiency evaluation of interpreters.
5.3.2 Comparison with the State-of-the-art Interpreters.
Fig. 4, Fig. 5, and Fig. 6 illustrate the evaluation of various interpreters across the ISCXTor2016, CIC-DoHBrw-2020, and UNSW-NB15 datasets, highlighting performance differences across three critical dimensions: fidelity, stability, and robustness. Specifically, our interpreter achieves the highest fidelity on the ISCXTor2016 and UNSW-NB15 datasets. Although some interpreters achieve higher fidelity on the CIC-DoHBrw-2020 dataset, their fidelity is lower than ours on other datasets. A higher fidelity suggests that our method effectively identifies critical features that have a substantial influence on the model’s decision-making process. In terms of stability, our interpreter consistently produces similar results across multiple runs, ensuring reliable interpretations. Additionally, our method demonstrates strong robustness, maintaining accurate interpretations even when noise is introduced to the data. Overall, compared to other state-of-the-art interpreters, our model provides stable and reliable decision-making with higher credibility. Besides, Fig. 7, in turn, presents the results of the efficiency evaluation on all three datasets. KAM demonstrates a superior balance across all these dimensions compared to other methods on all the evaluation datasets.
@@
As shown in Fig. 4, we observe that in the ISCXTor2016 dataset, KAM achieves the highest fidelity compared to LIME, DeepLift, DiCE, CADE, DeepSHAP, SHAP, xNIDS, and Anchors, while the stability of KAM is higher than that of LIME, DiCE, CADE, SHAP, xNIDS, and Anchors. Additionally, KAM’s robustness is higher than that of LIME, DeepLift, DiCE, CADE, DeepSHAP, SHAP, xNIDS, and Anchors. Although the robustness of COIN is higher than KAM on the ISCXTor2016 dataset, its stability is slightly worse than KAM, and its efficiency is significantly lower, being nearly 100 times slower than KAM. Notably, the fidelity of all interpreters is relatively low on ISCXTor2016 because anomaly samples, such as Tor traffic, are more susceptible to label flipping, as modifications to their key features are more likely to alter model predictions. However, the dataset’s significant class imbalance—reflected in a normal-to-anomalous ratio of approximately 4.85:1—leads to a lower overall LFR. In contrast, as depicted in Fig. 5 and Fig. 6, the CIC-DoHBrw-2020 and UNSW-NB15 dataset has a normal-to-anomalous ratio of about 3.31:1 and 0.56:1, offering a relatively more balanced scenario for evaluating fidelity.
As presented in Fig. 5, KAM’s robustness is optimal compared with all the baselines on the CIC-DoHBrw-2020 dataset. Besides, the stability of KAM outperforms LIME, DiCE, CADE, SHAP, COIN, xNIDS, and Anchors, and its fidelity is also higher compared to DeepLift, CADE, DeepSHAP, SHAP, COIN, and xNIDS. Despite that, LIME, DiCE, and ANchors obtain a higher fidelity than KAM; their stability is way worse compared to KAM. Although the stability of DeepLift is similar to that of KAM, its robustness is lower than that of KAM. KAM and DeepSHAP share a similar stability while KAM’s fidelity is much better than DeepSHAP.
Fig. 6 demonstrates the evaluation results of KAM compared with the state-of-the-art interpreters regarding fidelity, stability, and robustness on the UNSW-NB15 dataset. The results show that KAM has the best performance regarding fidelity, robustness, and stability compared to all the state-of-the-art interpreters. The stability of DeepSHAP and DeepLift is similar to KAM; however, their fidelity and robustness are much lower than KAM.
Moreover, according to Fig. 7, we can see that KAM is more efficient than SHAP, COIN, DiCE, CADE, and Anchors on the ISCXTor2016 dataset. The efficiency of LIME is similar to that of KAM, while its stability, fidelity, and robustness are worse than KAM on ISCXTor2016. DeepLift, xNIDS, and DeepSHAP show better efficiency than KAM; however, their fidelity and robustness are much lower than KAM on ISCXTor2016. In addition, on the CIC-DoHBrw-2020 dataset, the efficiency of KAM is much higher than SHAP, COIN, DiCE, CADE, Anchors, and LIME, and KAM’s efficiency is better than COIN, DiCE, CADE, Anchors, and LIME on the UNSW-NB15 dataset. KAM’s efficiency is not as good as that of DeepLift, xNIDS, and DeepSHAP on both CIC-DoHBrw-2020 and UNSW-NB15 datasets due to the two-branch design of KAM (i.e., with both global and local interpreters for higher reliable interpretations), but it has a higher fidelity and robustness on CIC-DoHBrw-2020 and UNSW-NB15. Even though SHAP has similar efficiency compared to KAM on the UNSW-NB15 dataset, its fidelity and robustness are way worse than KAM.
5.3.3 Comparison with the State-of-the-art Semi-supervised Models.
@@
As shown in Table 5, Table 6, and Table 7, ToS outperforms other methods on the ISCXTor2016, CIC-DoHBrw-2020, and UNSW-NB15 datasets. Overall, ToS achieves SPAUC as high as 67.42%, 55.94%, and 57.02% on ISCXTor2016, CIC-DoHBrw-2020, and UNSW-NB15 datasets with about 1.33%, 1.53%, and 3.81% improvement, respectively. The accuracy, precision, recall, F1-score, and SPAUC of ToS are the best on the ISCXTor2016, CIC-DoHBrw-2020, and UNSW-NB15 datasets compared with the state-of-the-art semi-supervised models (i.e., instantT, ACR, Clustering*, LVM, M3S, and CETP). Although ToS’s FAR is slightly worse than Clustering*, its accuracy, precision, recall, F1-score, MDR, and SPAUC are 6.38%, 35.69%, 22.23%, 29.95%, 45.09%, and 17.44% better than those of Clustering*. Similarly, instantT’s MDR is 0.58% lower than ToS, but its accuracy, precision, recall, F1-score, FAR, and SPAUC are 5.13%, 14.54%, 2.83%, 6.36%, 6.24%, and 11.59% worse than ToS on the ISCXTor2016 dataset, as illustrated in Table 5.
As depicted in Table 6, a similar phenomenon is observed for the CIC-DoHBrw-2020 dataset. In particular, the FAR of instantT and the MDR of ACR are better than ToS, but they do not perform as well as ToS on other evaluation metrics on the CIC-DoHBrw-2020 dataset. Similarly, ToS outperforms the state-of-the-art semi-supervised models in terms of accuracy, precision, recall, F1-score, FAR, and SPAUC. Despite the fact that instantT’s FAR and ACR’s MDR are lower than those of ToS, ToS significantly outperforms them in terms of other metrics on the CIC-DoHBrw-2020 dataset.
As presented in Table 7, the comparison results on the UNSW-NB15 dataset demonstrate that ToS archives the best SPAUC and F1-score for detecting anomalies from browser-based DoH activities compared with the state-of-the-art semi-supervised models. In particular, ToS’s SPAUC is 5.58%, 4.29%, 4.68%, 5.52%, 5.40%, and 4.46% higher than that of instantT, ACR, Clustering*, LVM, M3S, and CETP on the UNSW-NB15 dataset. Although the accuracy, precision, and MDR of ToS are worse than instantT, its FAR and SPAUC are 36.12% and 5.58% better than those of instantT.
@@
Additionally, as illustrated in Fig. 8 and Fig. 9, ToS strikes the optimal balance between F1-score, SPAUC, FAR, and MDR. Specifically, ToS is optimal for anomaly detection on ISCXTor2016, CIC-DoHBrw-2020, and UNSW-NB15 in terms of F1-score and SPAUC compared with all the state-of-the-art approaches, as presented in Fig. 8. As illustrated in Fig. 9, ToS outperforms the comparative state-of-the-art semi-supervised algorithms regarding FAR and MDR on the ISCXTor2016 dataset, and it has the lowest FAR on the UNSW-NB15 dataset. Although the instantT method exhibits an exceptionally low FAR on the CIC-DoHBrw-2020 dataset, its MDR is extremely high, nearing 80%. Similarly, while ACR demonstrates superior MDR performance compared to ToS on the CIC-DoHBrw-2020 dataset, both suffer from significantly elevated FAR values.
@@
Table 5: Performance (%) of ToS compared with the state-of-the-art semi-supervised models on ISCXTor2016.
@@
Table 6: Comparison with the state-of-the-art semi-supervised models’ performance (%) on CIC-DoHBrw-2020.
@@
Table 7: Comparison with the state-of-the-art semi-supervised models’ performance (%) on UNSW-NB15.
Model Acc. Pre. Rec. F1. FAR MDR SPAUC
instantT 78.74 85.72 76.39 76.48 46.85 0.35 51.44
ACR 78.20 78.02 77.81 77.89 26.03 20.64 52.73
Clustering* 71.92 72.05 72.26 71.88 24.37 31.09 52.34
LVM 78.10 82.78 76.03 76.21 44.46 3.47 51.50
M3S 76.33 77.76 74.89 75.15 39.36 10.85 51.62
CETP 71.54 72.01 72.13 71.54 22.11 33.63 52.56
ToS (ours) 78.38 79.38 78.38 78.38 10.73 30.49 57.02
Figure 8: Performance of ToS compared with the state-of-the-art semi-supervised models: F1-score vs. SPAUC.
Figure 9: Performance of ToS compared with the state-of-the-art semi-supervised models: FAR vs. MDR.
6 Related Work